# Reconfigurable Intelligent Surfaces for 6G Mobile Networks: An Industry R&D Perspective

### v1

@@ Abstract
The reconfigurable intelligent surface (RIS) technology is a potential solution to enhance network capacity and coverage without significant investiment into additional infrastructure in 6G networks. This work highlights the interest of the mobile communication industry on RIS, and discusses the development of liquid crystal-based RIS for improved energy efficiency and coverage in the millimeter-wave band. Furthermore, the paper discusses perspectives and insights from an industry R&D point of view, addressing relevant use cases, technical requirements, implementation challenges, and practical considerations for RIS deployment optimization in the context of 6G networks.
@@ I Introduction
The fifth-generation mobile networks are currently being rolled out worldwide. Theoretically, 5G should provide data rates of up to 10 Gbps in average and latencies below 10 ms. However, it is clear that with future applications, for example in connection with telepresence, mixed realities or for highly specialized use in an industrial context, the re-quirements for data rate, reliability, connection density and latency will continue to increase. Among other targets, 6G aims at larger transmission rates, lower latencies, and more energy efficiency. It provides connectivity even under unfavorable conditions in order to enable safety-critical applications such as the connection for unmanned autonomous vehicles. These targets cannot be achieved by simply deploying additional base stations, but requires alternative, innovative solutions. With regard to the long-term national and international climate goals as well as the social acceptance of 6G, it is crucial to focus on energy efficiency in all necessary developments and network expansions. Sustainability is highlighted as an essential criterion in which 6G should build upon 5G.
@@
Because of its potential to significantly improve the performance of radio access network, RISs have caught the attention of academia, industry, funding bodies, and regulators, and, therefore, significant investments have been directed towards the development and evaluation of this technology towards 6G. For example, the 6G LICRIS (Liquid Crystal Reconfigurable Intelligent Surfaces) project aims to develop RIS based on liquid crystal (LC) for surface reconfigurability from a holistic network perspective. Compared to semiconductor-based RIS, LC is more energy efficient in higher frequencies, and is a promising technology candidate for RIS hardware implementation.
@@ II-A Relevant Use Cases
Use cases for RIS are currently discussed in the radio communication industry, starting from coverage extension towards smart antenna applications. However, standardisation bodies, such as 3GPP or ETSI have just started to analyse possible opportunities. One of the first SDO working group to look into RIS specifications and applications has been established in ETSI early in 2021, namely the ETSI ISG RIS. Still, the technical group reports are not published and in a drafting stage. So far, they have identified the following potential use cases for RIS deployments:
@@
    Signal enhancement for blind users: the direct LOS is obstructed by a building. This reduces the signal quality on receiver side massively. By using the nearby installed RIS, a “quasi-LOS” connection can be established and the signal quality is significantly improved. This application is particularly important in high-frequency scenarios (mmWave and THz communication), as these kind of signals are highly sus-ceptible to blockage and pathloss [8].
@@
    Interference suppression: By deploying RISs on cell edges the desired signal strengths can be improved compared to the inter-cell interference (interference by other neigh-boring base stations/cells) signal strengths. This is done by passive beamforming as the signals are fed to the target direction only. Following this the spread of signals into areas that should not actually be covered is reduced [9].
@@
    Improve channel rank conditions: RISs allow to manipulate the wireless channel by adding additional propagation paths into the desired direction. By this the channel rank condi-tion can be improved [9].
@@
    Mounting of RIS on unmanned aerial vehicle (UAV) To assist and enhance terrestrial communication, RISs can be attached to UAVs [12]. This idea is similar to the mounting of RIS modules to buildings. However, attachment to UAVs provides even more degrees of freedom with regard to positioning. It extends the coverage of the wireless sig-nals by beamforming with UAV trajectory optimization [9].
@@
    RIS to support simultaneous wireless information and power transfer (SWIPT): Internet of things (IoT) devices like small sensors that support our todays life are more and more spread into the world. Traditionally these devices are equipped with small batteries. One emerg-ing approach studies the power supply of these devices by using the radiated energy for the wire-less information transfer (Simultaneous wireless information and power transfer). By introducing RISs in such scenarios the energy collecting and data transmission capabilities can be boosted [9].
@@
    RIS assisted non-orthogonal multiple access (NOMA): Compared to conventional multiple input multiple output (MIMO)-NOMA, RIS assisted NOMA can improve the spectral efficiency and user connectivity by overcoming the challenges of random channel fluctuations, blockage and user mobility [14].
@@
    Enhancing indoor signal strength: Due to the intrinsic impedance difference between free-space and glass, the signal magnitude through windows is reduced drastically. To overcome this challenge, transparent reflecting intelligent surfaces, also known as transparent intelligent surfaces, represent a suitable solution for improving the signal quality without greatly impairing the actual function of a glass surface, leading to an optical transparency of 80% [15].
@@ II-B1 Regulation
As RIS modules and RIS based products will be primarily deployed as radio network components to extend coverage, as depicted in Figure 1, or offer radio beam shaping capabilities to transmitters, they obviously have to meet regulatory requirements according to RED. Thus, it is expected that there will be the need for RIS specific harmonized requirements standardization in the Euro-pean Union. Still, there is no request by the European Commission, however, standardization bod-ies like ETSI have already started into this, namely the ETSI ISG RIS (ETSI industry specification Group for RIS).
@@ II-B2 Conformance
As soon as RIS will be used to manipulate the impinging signal with respect to radio parameters, e.g., beamforming, and signal quality parameters,e.g., modulation quality, there is a need to test those capabilities according to the standard. So far, as explained before, standard developing organizations (SDOs) are still in an early phase regarding RIS specifications, and the potential need for conformance and performance assessment is yet to be specified.
@@ II-B3 Operational Requirements Gaps
Possibly the most serious concern of the industry regarding the operation of RIS is the fact that it will not be irradiated by wanted signals only, but also any other radio signal out there. Thus, co-existence issues need to be addressed with great interest. According to regulations, radio equipment shall not harm other services, thus, co-existence can be regarded as a regulatory issue, too. Therefore, co-existence testing scenarios need to be considered, including inband (wanted signal spectrum) and out of band, e.g., adjacent spectrum signals. Nevertheless, this is obviously a wide area and needs to be addressed by dedicated investigations, which is out of the scope of this work.
@@ II-C2 Regulatory and Multi-Operator Interference Issues
One of the major challenges for RIS deployment in multi-operator scenarios is managing interference among different network operators, ensuring fairness in resource allocationallocation, and avoiding destructive interference. As RISs can potentially modify the propagation environment across multiple operators, the allocation of resources and the concurrent usage of multiple RISs deployed by different operators become critical concerns. The wideband properties and non-linear saturation effects of RISs introduce challenges related to unwanted interference and regulatory compliance [20]. When an RIS is deployed and controlled by one operator, it can cause unpredictable changes in the propagation environment for other operators in the nearby spectrum. Such alterations in the propagation environment can lead to issues such as additional reflections affecting cell coverage, neighbour-cell interference, and may disturb network planning. Furthermore, rapid variations in reflection patterns can disturb instantaneous spatial multiplexing and link adaptation for both reciprocity-based and downlink CSI-based scenarios. Research in this area is in its nascent stages, but preliminary work suggests the need for collaborative efforts among operators. Addressing these challenges necessitates the development of interference mitigation techniques and careful consideration of regulatory guidelines.
@@ II-C3 Competitive Technologies
The benefits of RISs must also be assessed compared to competitive technologies like the decode-and-forward relay network-controlled repeaters (NCRs) [24, 25] or the network-controlled repeater (NCRs) [26]. While these solutions may be more costly alternatives compared to RIS, they offer better controllability and potentially larger coverage, as RISs are more susceptible to multiplicative fading. The path loss of the BS-RIS/NCR-UE link is a product of the individual link path losses. It can be compensated for due to the amplification capabilities in the NCR case. For the case of RIS, it requires a close deployment of the RIS to either the BS or the UE to enable a signal improvement or a very large RIS with an extensive element count. Besides the production costs, deployment issues and maintenance costs, which all increase with size, a very large RIS leads to problems described in the earlier paragraph on channel estimation challenges. For a fair comparison between RIS and NCRs all named advantages and disadvantages must be added up to a total cost of ownership per deployment. To evaluate the economic choice between RIS and NCR, the network operators will have to consider the cost for coverage, which is a sum of the deployment costs per area. Because several cheaper RISs are required to cover the same area as one NCR, as stated earlier, the lower RIS deployment cost must level out this factor. The many aspects of the cost for coverage lead to a complex calculation. This evaluation is seen as a crucial aspect of future research. A more extensive comparison of RISs and NCRs is provided in [7].
@@ III Hardware Design
RISs are based on unit cell elements that apply a phase shift to the incoming waves and re-radiate them back to the cell environment. In this section, hardware concepts that are under development in the 6G LICRIS project are briefly discussed to illustrate novel approaches to RIS design. In particular, the proposed concepts follow a passive design, reflecting the outgoing RF wave and saving power compared to base station antennas, with the goal of enhancing coverage, spectral efficiency, and beam management while improving energy efficiency. Two different RIS models are proposed at the 28 GHz band, which could be tested in indoor scenarios.
@@
The first model consists of a multi-resonant RIS integrated in liquid crystal. This topology consists a layer of liquid crystal comprised between two layers of glass substrate. Different metallic resonant structures are printed over the glass making use of photolitography, this way different resonant structures will lay over and under the liquid crystal, creating a multi-resonant topology whose resonant frequencies depends on the pollarization of the molescules of the LC-material. When multiple resonances are used, a wide bandiwdth is expected, in this case a bandwidth of around 6 GHz with a central frequency of 28 GHz is expected. This large bandwidth will enable the use of a high data-rate. This topology is compatible with a single pollarization and a frontview of what the RIS would look like and the expected phase of a single unit cell of the RIS topology are shown in Figure 3. The phase different works between 24 and 32 GHz approximately with a difference of around 300, which is relatively constant in the whole frequency range.
(a) Stackup view of an unit cell of the stacked-patch RIS
@@
The second type of RIS consists of a coupled patch RIS. In this topology the radiating element is a patch which is coupled to an reflective LC phaseshifter. One the wave reaches the patch, it is coupled into the active phaseshifter throught an slot in the ground plane. This phaseshifter provides a changeable phaseshift which is between 0 and 360 a bandwidth of around 3 GHz is expected and it is compatible with two orthogonal polarizations. The whole stack-up of the RIS and the phaseshifter for different values of permittivity are shown in Figure 4.
@@ IV-A Propagation Modeling
Semi-empirical and/or geometric RF propagation models offer a good compromise. The measures that need to be taken into account are firstt of all the received power, in a more sophisticated model, beyond the scope of the current work, also the impact of the MIMO path through the RIS are of interest. Since RIS can be most efficiently used in higher frequency beyond 6 GHz, RF propagation in scenarios where RIS are used is typically limited by line-of-sight (LoS) between transmitter and receiver. It is the intention of using RIS to supplement the LoS-pathes by reflections on RIS. While semi-empirical models are widely used in sub-6GHz-mobile radio scenarios (or historically sub-2GHz), geometric models are less limited in its frequency application ranges, since they are based on deterministic approaches.
@@
Semi-empirical models offer the advantage of relatively low computational complexity compared to other RF propagation modeling approaches such as ray tracing and full-wave simulations, which can be computationally intensive and time consuming, and also have limitations. They are based on empirical measurements that may not always capture the full complexity of real-world environments, and they may not accurately predict the effects of new technologies and deployment scenscenarios that differ from the assumptions of the original model. Therefore, careful validation and calibration of semi-empirical models using real-world measurements and data is essential to ensure their accuracy and reliability.
@@
Geometric models are based on the concept that RF signals propagate in straight lines and their behavior can be predicted by considering the geometry of the environment in which they propagate. These models use parameters such as the distance between the transmitter and receiver, the height of the antennas, and the surrounding terrain to calculate the path loss of RF signals. For example, the two-beam ground reflection model accounts for reflections of RF signals from the ground by assuming that the signal travels in two paths: a direct path from the transmitter to the receiver and another reflected path from the ground to the receiver. The model takes into account the heights of the antennas and the distance between the transmitter and receiver to calculate the path loss. The two-beam ground reflection model is commonly used to predict RF propagation in urban environments where ground reflections can significantly affect the path loss. A similar approach can be considered for RF prediction with RIS, assuming that the reflection point is from the RIS rather than the ground.
@@
Knife-edge models are named for the sharp edge of the obstacle, which is considered to be the most significant factor affecting the diffraction of RF signals. The shape, height and distance of the obstacle from the transmitter and receiver are key parameters in these models. One of the most commonly used knife-edge models is the single knife-edge diffraction model, which assumes that the RF signal encounters a single sharp edge along its propagation path. The model calculates the path loss based on the height and distance of the obstacle, the height of the antennas, the frequency of the RF signal, and the wavelength of the signal.
For the integration of RIS into RF propagation models, diffraction is certainly of less interest than reflection. However, both modeling approaches could be combined by a combined parameter estimation of their parameters with methods of mathematical optimization, including machine learning/artificial intelligence approaches. Such an approach depends strongly on the availability of suitable training data. For RIS such data are currently hardly available. A combination of simulation and ray tracing may provide artificial training data, similar to the approach of RadioUNet [28].

### v2

@@ Abstract
The reconfigurable intelligent surface (RIS) technology is a potential solution to enhance network capacity and coverage without significant investment into additional infrastructure in 6G networks. This work highlights the interest of the mobile communication industry on RIS, and discusses the development of liquid crystal-based RIS for improved energy efficiency and coverage in the millimeter-wave band. Furthermore, the paper discusses perspectives and insights from an industry R&D point of view, addressing relevant use cases, technical requirements, implementation challenges, and practical considerations for RIS deployment optimization in the context of 6G networks.
@@ I Introduction
The fifth-generation (5G) mobile networks are currently being rolled out worldwide. Theoretically, 5G should provide data rates of up to 10 Gbps in average and latencies below 10 ms. However, it is clear that with future applications, for example in connection with telepresence, mixed realities or for highly specialized use in an industrial context, the re-quirements for data rate, reliability, connection density and latency will continue to increase. Among other targets, 6G aims at larger transmission rates, lower latencies, and more energy efficiency. It provides connectivity even under unfavorable conditions in order to enable safety-critical applications such as the connection for unmanned autonomous vehicles. These targets cannot be achieved by simply deploying additional base stations, but requires alternative, innovative solutions. With regard to the long-term national and international climate goals as well as the social acceptance of 6G, it is crucial to focus on energy efficiency in all necessary developments and network expansions. Sustainability is highlighted as an essential criterion in which 6G should build upon 5G.
@@
Because of its potential to significantly improve the performance of radio access network, RISs have caught the attention of academia, industry, funding bodies, and regulators, and, therefore, significant investments have been directed towards the development and evaluation of this technology towards 6G. For example, the 6G LICRIS (Liquid Crystal Reconfigurable Intelligent Surfaces) project aims to develop RIS based on liquid crystal (LC) for surface reconfigurability from a holistic network perspective. Compared to semiconductor-based RIS, LC is more energy efficient in higher frequencies, it is scalable, cost-effective, and a promising technology candidate for RIS hardware implementation.
@@ II-A Relevant Use Cases
Use cases for RIS are currently discussed in the radio communication industry, starting from coverage extension towards smart antenna applications. However, standardisation bodies, such as 3GPP or ETSI have just started to analyse possible opportunities. One of the first standard developing organizations (SDO) working group to look into RIS specifications and applications has been established in ETSI early in 2021, namely the ETSI ISG RIS (ETSI industry specification group for RIS). Still, the technical group reports are not published and in a drafting stage. So far, they have identified the following potential use cases for RIS deployments:
@@
    Signal enhancement for blind users: the direct LOS is obstructed by a building. This reduces the signal quality on receiver side massively. By using the nearby installed RIS, a “quasi-LOS” connection can be established and the signal quality is significantly improved. This application is particularly important in high-frequency scenarios (mmWave and THz communication), as these kind of signals are highly susceptible to blockage and pathloss [8].
@@
    Interference suppression: By deploying RISs on cell edges the desired signal strengths can be improved compared to the inter-cell interference (interference by other neighboring base stations/cells) signal strengths. This is done by passive beamforming as the signals are fed to the target direction only. Following this the spread of signals into areas that should not actually be covered is reduced [9].
@@
    Improve channel rank conditions: RISs allow to manipulate the wireless channel by adding additional propagation paths into the desired direction. By this the channel rank condition can be improved [9].
@@
    Mounting of RIS on unmanned aerial vehicle (UAV) To assist and enhance terrestrial communication, RISs can be attached to UAVs [12]. This idea is similar to the mounting of RIS modules to buildings. However, attachment to UAVs provides even more degrees of freedom with regard to positioning. It extends the coverage of the wireless signals by beamforming with UAV trajectory optimization [9].
@@
    RIS to support simultaneous wireless information and power transfer (SWIPT): Internet of things (IoT) devices like small sensors that support our todays life are more and more spread into the world. Traditionally these devices are equipped with small batteries. One emerging approach studies the power supply of these devices by using the radiated energy for the wireless information transfer (SWIPT). By introducing RISs in such scenarios the energy collecting and data transmission capabilities can be boosted [9].
@@
    RIS assisted non-orthogonal multiple access (NOMA): Compared to conventional multiple input multiple output (MIMO)-NOMA, RIS assisted NOMA can improve the throughout, spectral efficiency, and user connectivity by overcoming the challenges of random channel fluctuations, blockage and user mobility [14].
@@
    Enhancing indoor signal strength: Due to the intrinsic impedance difference between free-space and glass, the signal magnitude through windows is reduced drastically. To overcome this challenge, transparent reflecting intelligent surfaces, also known as transparent intelligent surfaces, represent a suitable solution for improving the signal quality without greatly impairing the actual function of the glass, leading to an optical transparency of 80% [15].
@@ II-B1 Regulation
As RIS modules and RIS based products will be primarily deployed as radio network components to extend coverage, as depicted in Figure 1, or offer radio beam shaping capabilities to transmitters, they obviously have to meet regulatory requirements according to RED. Thus, it is expected that there will be the need for RIS specific harmonized requirements standardization in the European Union. Still, there is no request by the European Commission, however, standardization bod-ies like ETSI have already started into this, namely the ETSI ISG RIS.
@@ II-B2 Conformance
As soon as RIS will be used to manipulate the impinging signal with respect to radio parameters, e.g., beamforming, and signal quality parameters,e.g., modulation quality, there is a need to test those capabilities according to the standard. So far, as explained before, SDOs are still in an early phase regarding RIS specifications, and the potential need for conformance and performance assessment is yet to be specified.
@@ II-B3 Operational Requirements Gaps
Possibly the most serious concern of the industry regarding the operation of RIS is the fact that it will not be irradiated by wanted signals only, but also any other radio signal out there. Thus, co-existence issues need to be addressed with great interest. According to regulations, radio equipment shall not harm other services, thus, co-existence can be regarded as a regulatory issue, too. Therefore, coexistence testing scenarios need to be considered, including in-band (wanted signal spectrum) and out of band, e.g., adjacent spectrum signals. Nevertheless, this is obviously a wide area and needs to be addressed by dedicated investigations, which is out of the scope of this work.
@@ II-C2 Regulatory and Multi-Operator Interference Issues
One of the major challenges for RIS deployment in multi-operator scenarios is managing interference among different network operators, ensuring fairness in resource allocation, and avoiding destructive interference. As RISs can potentially modify the propagation environment across multiple operators, the allocation of resources and the concurrent usage of multiple RISs deployed by different operators become critical concerns. The wideband properties and non-linear saturation effects of RISs introduce challenges related to unwanted interference and regulatory compliance [20]. When an RIS is deployed and controlled by one operator, it can cause unpredictable changes in the propagation environment for other operators in the nearby spectrum. Such alterations in the propagation environment can lead to issues such as additional reflections affecting cell coverage, neighbour-cell interference, and may disturb network planning. Furthermore, rapid variations in reflection patterns can disturb instantaneous spatial multiplexing and link adaptation for both reciprocity-based and downlink CSI-based scenarios. Research in this area is in its nascent stages, but preliminary work suggests the need for collaborative efforts among operators. Addressing these challenges necessitates the development of interference mitigation techniques and careful consideration of regulatory guidelines.
@@ II-C3 Competitive Technologies
The benefits of RISs must also be assessed compared to competitive technologies like the decode-and-forward relay [24, 25] or the network-controlled repeater (NCRs) [26]. While these solutions may be more costly alternatives compared to RIS, they offer better controllability and potentially larger coverage, as RISs are more susceptible to multiplicative fading. The path loss of the BS-RIS/NCR-UE link is a product of the individual link path losses. It can be compensated by amplification capabilities in the NCR case. For the case of RIS, it requires a close deployment of the RIS to either the BS or the UE to enable a signal improvement or a very large RIS with an extensive element count. Besides the production costs, deployment issues and maintenance costs, which all increase with size, a very large RIS leads to problems described in the earlier paragraph on channel estimation challenges. For a fair comparison between RIS and NCRs all named advantages and disadvantages must be added up to a total cost of ownership per deployment. To evaluate the economic choice between RIS and NCR, the network operators will have to consider the cost for coverage, which is a sum of the deployment costs per area. Because several cheaper RISs are required to cover the same area as one NCR, as stated earlier, the lower RIS deployment cost must level out this factor. The many aspects of the cost for coverage lead to a complex calculation. This evaluation is seen as a crucial aspect of future research. A more extensive comparison of RISs and NCRs is provided in [7].
@@ III Hardware Design
RISs are based on unit cell elements that apply a phase shift to the incoming waves and reradiate them back to the cell environment. In this section, hardware concepts that are under development in the 6G LICRIS project are briefly discussed to illustrate novel approaches to RIS design. In particular, the proposed concepts follow a passive design, reflecting the outgoing RF wave and saving power compared to base station antennas, with the goal of enhancing coverage, spectral efficiency, and beam management while improving energy efficiency. Two different RIS models are proposed at the 28 GHz band, which could be tested in indoor scenarios.
@@
The first model consists of a multi-resonant RIS integrated in liquid crystal. This topology consists of a layer of liquid crystal comprised between two layers of glass substrate. Different metallic resonant structures are printed on the glass making use of photolitography. This way different resonant structures are on top of and below the liquid crystal layer, creating a multi-resonant topology whose resonant frequencies depend on the polarization of the molecules of the LC-material. When multiple resonances are used, a wide bandwidth is expected, in this case a bandwidth of around 6 GHz with a central frequency of 28 GHz is expected. This large bandwidth will enable the use of a high data-rate. This topology is compatible with a single polarization. A frontview of how the RIS would look like and the expected phase of a single unit cell of the RIS topology are shown in Figure 3. The phase different works between 24 and 32 GHz approximately with a difference of around 300, which is relatively constant in the whole frequency range.
(a) Stackup view of a unit cell of the stacked-patch RIS
@@
The second type of RIS consists of a coupled patch RIS. In this topology the radiating element is a patch which is coupled to a reflective LC phaseshifter. Once the wave reaches the patch, it is coupled into the active phaseshifter through a slot in the ground plane. This phaseshifter provides a changeable phaseshift which is between 0 and 360. A bandwidth of around 3 GHz is expected and it is compatible with two orthogonal polarizations. The whole stack-up of the RIS and the phaseshifter for different values of permittivity are shown in Figure 4.
@@ IV-A Propagation Modeling
Semi-empirical and/or geometric RF propagation models offer a good compromise. The measures that need to be taken into account are first of all the received power (see Figure 5), in a more sophisticated model, beyond the scope of the current work, also the impact of the MIMO path through the RIS are of interest. Since RIS can be most efficiently used in higher frequency beyond 6 GHz, RF propagation in scenarios where RIS are used is typically limited by line-of-sight (LoS) between transmitter and receiver. It is the intention of using RIS to supplement the LoS-pathes by reflections on RIS. While semi-empirical models are widely used in sub-6GHz-mobile radio scenarios (or historically sub-2GHz), geometric models are less limited in its frequency application ranges, since they are based on deterministic approaches.
@@
Semi-empirical models offer the advantage of relatively low computational complexity compared to other RF propagation modeling approaches such as ray tracing and full-wave simulations, which can be computationally intensive and time-consuming, and also have limitations. They are based on empirical measurements that may not always capture the full complexity of real-world environments, and they may not accurately predict the effects of new technologies and deployment scenscenarios that differ from the assumptions of the original model. Therefore, careful validation and calibration of semi-empirical models using real-world measurements and data is essential to ensure their accuracy and reliability.
@@
Geometric models are based on the concept that RF signals propagate in straight lines and their behavior can be predicted by considering the geometry of the environment in which they propagate (see Figure 6, right-hand side). These models use parameters such as the distance between the transmitter and receiver, the height of the antennas, and the surrounding terrain to calculate the path loss of RF signals. For example, the two-beam ground reflection model accounts for reflections of RF signals from the ground by assuming that the signal travels in two paths: a direct path from the transmitter to the receiver and another reflected path from the ground to the receiver. The model takes into account the heights of the antennas and the distance between the transmitter and receiver to calculate the path loss. The two-beam ground reflection model is commonly used to predict RF propagation in urban environments where ground reflections can significantly affect the path loss. A similar approach can be considered for RF prediction with RIS, assuming that the reflection point is from the RIS rather than the ground.
@@
Knife-edge models are named for the sharp edge of the obstacle, which is considered to be the most significant factor affecting the diffraction of RF signals (see Figure 6, left-hand side). The shape, height and distance of the obstacle from the transmitter and receiver are key parameters in these models. One of the most commonly used knife-edge models is the single knife-edge diffraction model, which assumes that the RF signal encounters a single sharp edge along its propagation path. The model calculates the path loss based on the height and distance of the obstacle, the height of the antennas, the frequency of the RF signal, and the wavelength of the signal.
For the integration of RIS into RF propagation models, diffraction is certainly of less interest than reflection. However, both modeling approaches could be combined by a combined parameter estimation of their parameters with methods of mathematical optimization, including machine learning/artificial intelligence approaches (see Figure 7). Such an approach depends strongly on the availability of suitable training data. For RIS such data are currently hardly available. A combination of simulation and ray tracing may provide artificial training data, similar to the approach of RadioUNet [28].